### v2

# Are KANs Effective for Multivariate Time Series Forecasting?

Abstract
Multivariate time series forecasting is a crucial task that predicts the future states based on historical inputs. Related techniques have been developing in parallel with the machine learning community, from early statistical learning methods to current deep learning methods. Despite their significant advancements, existing methods continue to struggle with the challenge of inadequate interpretability. The rise of the Kolmogorov-Arnold Network (KAN) provides a new perspective to solve this challenge, but current work has not yet concluded whether KAN is effective in time series forecasting tasks. In this paper, we aim to evaluate the effectiveness of KANs in time-series forecasting from the perspectives of performance, integrability, efficiency, and interpretability. To this end, we propose the Multi-layer Mixture-of-KAN network (MMK), which achieves excellent performance while retaining KAN’s ability to be transformed into a combination of symbolic functions. The core module of MMK is the mixture-of-KAN layer, which uses a mixture-of-experts structure to assign variables to best-matched KAN experts. Then, we explore some useful experimental strategies to deal with the issues in the training stage. Finally, we compare MMK and various baselines on seven datasets. Extensive experimental and visualization results demonstrate that KANs are effective in multivariate time series forecasting. Code is available at: https://github.com/2448845600/EasyTSF.
1 Introduction
(a) Performance comparison
(b) Training Loss
Figure 1: Two experiental findings on the ETTh1 dataset. Subfigure (a) shows the univariate forecasting results of KANs with different base functions, where results are minmax-normalized and 0 means the minimum MSE in the row. Subfigure (b) reports the training loss of one/three layers of Linear/KAN models.
The multivariate time series forecasting (MTSF) task aims to predict future states of variables from input historical data. Related technologies have been applied to a broad scope of fields, such as financial investment, weather forecasting, and traffic management Gao et al. (2023); Savcisens et al. (2024); Han et al. (2024). Time series forecasting methods have long been inspired by advances in the machine learning community. The popularity of early statistical learning methods gave rise to SVR and ARMIA Holt (2004); recurrent neural networks (RNN), which are specially designed to process sequence data, are naturally applied to time series analysis Hewamalage et al. (2021); TFT Lim et al. (2021) and Informer Zhou et al. (2021) promote the application of transformer architecture in time series prediction; LSTF-Linear Zeng et al. (2023) rethinks the value of Linear. At present, various deep learning network architectures, such as RNN, CNN, Transformer, and MLP, are widely used in time series forecasting Wang et al. (2023); Zhou et al. (2024); Jia et al. (2024). The forecasting models derived from different network architectures have their own advantages in forecasting performance, running speed, and resource usage.
However, most popular time series forecasting models are based on deep learning-based modules, whose prediction process is blackbox and prediction results lack interpretability. These nontransparent methods are suspected to be unsuitable for tasks with low error tolerance, such as medicine, law, and finance. Some traditional, non-deep learning-based methods are interpretable, but it is difficult to achieve the expected performance. Fortunately, a novel kind of deep learning network, the Kolmogorov-Arnold Network (KAN) Liu et al. (2024b), which is based on the Kolmogorov-Arnold representation theorem (KART), has become a feasible solution to the above dilemma of choosing between performance and interpretability. KAN offers a pruning strategy that simplifies the trained KAN model into a set of symbolic functions, making it possible to analyze the mechanisms of the network, thereby significantly enhancing the interpretability.
Although KAN has the advantages of both performance and interpretability in theory, some works Yu et al. (2024a) have pointed out that it does not work well for real-world tasks. A few works Genet and Inzirillo (2024); Vaca-Rubio et al. (2024) have tried to apply KAN to time series analysis, but they have not used the characteristics of time series data to design efficient KAN-based networks, and lack fair comparisons on common datasets. Existing works cannot provide a convincing conclusion on whether KAN is effective in time series forecasting. Therefore, we aim to provide a reliable benchmark to comprehensively evaluate the effectiveness of the KAN in multivariate time series forecasting tasks for the first time. When applying KAN to time series forecasting, we encountered two challenges, as following.
One is how to select the most suitable KAN. The KAN proposed by Ziming Liu et al. is not only a model but also a general function-fitting framework. Subsequent studies SS (2024); Bozorgasl and Chen (2024); Aghaei (2024) use different base functions to replace the B-spline function for better performance on specific tasks. We report the experimental results about five KAN’s variants (KANs), each with a special base function, on seven univariates of the ETTh1 dataset, as illustrated in Figure 1(a). This makes it cumbersome to determine the appropriate KAN for achieving the best performance. The other is how to train the deeper KAN. We build a single-layer KAN model that can achieve similar performance with a single-layer Linear model. But as the layer of the KAN model increases, the performance drops significantly. We visualize the values of loss in 10 times repeated trainings in Figure 1(b) and find that the initial loss of KAN is larger than that of Linear, and the loss of the 3-layer KAN model becomes significantly larger and more unstable.
To address the above challenges, we first apply the mixture-of-experts technique to KAN and obtain the Mixture-of-KAN (MoK) layer, which uses a gating network to adaptively assign the input to the matching KAN layers with special base functions according to the variable dimensions. This structure aggregates the advantages of KAN’s variants in different cases. Then, we build a simple yet efficient Multi-layer Mixture-of-KAN network (MMK) based on MoK. By using modules that do not compromise the ability to be converted into symbolic functions, MMK achieves excellent performance while retaining the interpretability from KAN. Next, we analyze the phenomenon of abnormal training loss and use some training strategies to improve the performance of the proposed MMK. Finally, we evaluate the effectiveness of KANs in multivariate time series forecasting from four perspectives: performance, integration, speed, and interpretability. We compare MMK with many popular time series forecasting models on seven commonly used datasets, and experimental results show that MMK achieves state-of-the-art performance in most cases. We integrate MoK into recently proposed models and find that it can improve their performance. And we visualize the output weights of the gating network in the MoK layer and analyze the connection between KAN’s symbolic function conversion ability and domain knowledge.
To sum up, the contributions of this paper include:
  • To the best of our knowledge, this is the first work that comprehensively discusses the effectiveness of the booming KANs for time series forecasting.
  • To validate our claims, we propose an interpretable Multi-layer Mixture-of-KAN network (MMK), which uses the mixture-of-KAN layer to aggregate the advantages of KAN’s variants. And we use some training strategies to improve performance.
  • We fairly compare MMK with many baselines on seven real-world datasets, and the experimental results show that MMK achieves the competitive performance.
In summary, we can conclude that KANs are effective in multivariate time series forecasting.
2 Related Work
2.1 Multivariate Time-Series Forecasting
Various network architectures (such as Transformer, CNN, and MLP) are competing in time series forecasting recently. Transformer-based time series forecasting models have strong performance but high time and memory complexity. Informer Zhou et al. (2021) proposed ProbSparse self-attention to reduce the complexity from $\mathcal{O}(T^{2})$ to $\mathcal{O}(T\log T)$. Pyraformer Liu et al. (2022b) utilizes the pyramid attention mechanism to capture hierarchical multi-scale time dependencies with a time and memory complexity of $\mathcal{O}(T)$. PatchTST Nie et al. (2023) and Crossformer Zhang and Yan (2023) use the patch operation to reduce the number of input tokens, thereby reducing time complexity. The performance of early Multi-layer perception-based (MLP-based) models is generally weaker than transformer-based methods. However, NLinear Zeng et al. (2023) and RLinear Li et al. (2023) combine different normalization methods with a single-layer MLP, and achieve performance that exceeds Transformer-based models on some datasets with extremely low computational cost. Recurrent neural networks (RNNs) are suitable for handling sequence data, making them a favorable choice for time series analysis. The size of the hidden state in RNN is independent of the input time series length, so recent RNN-based time series prediction models, such as SegRNN Lin et al. (2023) and WITRAN Jia et al. (2024), apply the RNN structure to time series forecasting with longer input. Convolutional neural networks (CNNs) are frequently used in time series forecasting models in the form of 1D convolution, such as ModernTCN donghao and wang xue (2024) and SCINet Liu et al. (2022a), but TimesNet Wu et al. (2023) takes a different approach by converting the 1D time series into a 2D matrix through the Fourier transform and then using 2D convolution for prediction.
2.2 Kolmogorov-Arnold Network
The Kolmogorov-Arnold representation theorem (KART) is the mathematical foundation of the Kolmogorov-Arnold Network (KAN) Liu et al. (2024b), which makes KAN more fitting and interpretable than Multi-Layer Perceptrons (MLP) based on the universal approximation theorem. Given an input tensor $\mathbf{x}\in\mathbb{R}^{n_{0}}$, the structure of $L$ layers in a KAN network can be represented as:
$\operatorname{KAN}(\mathbf{x})=\left(\boldsymbol{\Phi}_{L}\circ\boldsymbol{\Phi}_{L-1}\circ\cdots\circ\boldsymbol{\Phi}_{2}\circ\boldsymbol{\Phi}_{1}\right)\mathbf{x},$ (1)
where $\boldsymbol{\Phi}_{l},l\in[1,2,\cdots,L]$ is a KAN layer, and the output dimension of each KAN layer can be expressed as: $[n_{1},n_{2},\cdots,n_{L}]$. Therefore, the transform process of $j$-th feature in $l$-th layer can be formed as:
$\mathbf{x}_{l,j}=\sum_{i=1}^{n_{l-1}}\phi_{l-1,j,i}\left(\mathbf{x}_{l-1,i}\right),\quad j=1,\cdots,n_{l},$ (2)
where $\phi$ consists two parts: the spline function and the residual activation function with learnable parameters $w_{b},w_{s}$:
$\phi(x)=w_{a}\mathrm{SiLU}(x)+w_{b}\mathrm{Spline}(x),$ (3)
where $\mathrm{Spline}(\cdot)$ is a linear combination of B-spline functions $\mathrm{Spline}(x)=\sum_{i}c_{i}B_{i}(x)$.
Recent studies have proposed many variants of KAN to expand its application scope and improve its performance. Some work uses functions such as wavelet functions Bozorgasl and Chen (2024), Taylor polynomials Yu et al. (2024b) and Jacobi polynomials Aghaei (2024) instead of spline functions to improve the performance of KAN in different fields. We can unify these KANs using the generalized version of Equation 3 as:
$\phi(x)=w_{a}\mathbb{A}(x)+w_{b}\mathbb{F}(x).$ (4)
3 Multi-layer Mixture-of-KAN Network
3.1 Problem Defintion
In multivariate time series forecasting, given historical data $\mathcal{X}=[X_{1},\cdots,X_{T}]\in\mathbb{R}^{T\times C}$, where $T$ is the time steps of historical data and $C$ is the number of variates. The time series forecasting task is to predict $\mathcal{Y}=[X_{T+1},\cdots,X_{T+P}]\in\mathbb{R}^{P\times C}$ during future $P$ time steps.
3.2 Mixture-of-KAN Layer
Real-world time series data frequently exhibit non-stationarity, with their statistical properties (such as mean and variance) varying over time. Moreover, there are significant distribution discrepancies between variables in multivariate time series. This poses a significant challenge to time series forecasting techniques, inevitably impacting the KAN-based methods as well. Fortunately, when it comes to dealing with distribution shift, KAN has a unique characteristic compared to existing methods: it has many variants using different base functions. Considering that the special base function may be suitable for modeling certain data distributions, we try to combine several KANs into a single layer and adaptively schedule them according to the input data.
Following this idea, we propose the Mixture-of-KAN (MoK) layer, which is a combination of KAN and mixture-of-experts (MoE). The MoK layer uses a gating network to assign KAN layers to variables according to temporal features, where each expert is responsible for a specific part of the data. KAN and its variants only differ in the base function, so we use $\mathcal{K}(\cdot)$ to represent KANs uniformly in this paper. Our proposed MoK layer with $N$ experts can be simply formed as:
$\mathbf{x}_{l+1}=\sum_{i=1}^{N}\mathcal{G}(\mathbf{x}_{l})_{i}\mathcal{K}_{i}(\mathbf{x}_{l}),$ (5)
where $\mathcal{G}(\cdot)$ is a gating network. This mixture of expert structure can adapt to the diversity of time series, with each expert learning different parts of temporal features, thus improving performance on time series forecasting task.
The gating network is the key module of the MoK layer, which is responsible for learning the weight of each expert from the input data. The softmax gating network $\mathcal{G}_{\mathrm{softmax}}$ which uses a softmax function and a learnable weight matrix $\mathbf{w}_{g}$ to schedule input data as:
$\mathcal{G}_{\mathrm{softmax}}(\mathbf{x})=\mathrm{Softmax}(\mathbf{x}\mathbf{w}_{g}).$ (6)
This gating network is popular due to its simple structure. However, it activates all experts once, resulting in low efficiency when there are a large number of experts. Therefore, we adopt the sparse gating network Shazeer et al. (2017) which only activates the best matching top-$k$ experts. It adds Gaussian noise to input time series by $\mathbf{w}_{\mathrm{noise}}$, and uses $\mathrm{KeepTopK}$ operation to retain experts with the highest $k$ values. The sparse gating network can be described as:
$\mathcal{G}_{\mathrm{sparse}}(\mathbf{x})=\mathrm{Softmax}(\mathrm{KeepTopK}(H(\mathbf{x}),k)),$ (7)
$H(\mathbf{x})=\mathbf{x}\mathbf{w}_{g}+\mathrm{Norm}(\mathrm{Softplus}(\mathbf{x}\mathbf{w}_{\mathrm{noise}})),$ (8)
where $\mathrm{Norm}(\cdot)$ is standard normalization.
Figure 2: The architecture of MMK. Subfigure (a) is the structure of the mixture-of-KAN (MoK) layer; subfigure (b) is the pre-sample initialization strategy for the last KAN layer.
3.3 Network Architecture
The interpretability of KAN comes from the fact that the trained KAN model can be converted into a combination of many simple symbolic functions. In order to maintain this advantage, we use layers or operators with the same properties to build our KAN-based forecasting model.
RevIN
RevIN Kim et al. (2022) includes symmetrically structured normalization $\text{RevIN}^{+}$ and denormalization operations $\text{RevIN}^{-}$ in the variable dimension. Before sending data $X$ to the backbone network, $\text{RevIN}^{+}$ uses a learnable affine transformation to normalize the input time series of each variable. After the last MoK layer, outputs are denormalized to the original distribution space using the same affine transformation parameters in the first step by $\text{RevIN}^{-}$ operation. Each learnable parameter of RevIN corresponds to the statistical value of the time series variable, and it can be viewed as a symbolic function in form, so it does not harm the interpretability of the overall network.
MoK Block
Inspired by the ResNet He et al. (2016), we design a simple MoKBlock, which consists of a MoK layer, BatchNorm, and Dropout in sequence, and adds a residual structure. By stacking several MoKBlocks, MMK can obtain the stronger feature extraction capability.
$\operatorname{MoKBlock}(\boldsymbol{h})=\operatorname{Dropout}(\operatorname{BN}(\boldsymbol{h}+\operatorname{MoK}(\boldsymbol{h}))).$ (9)
Loss Function
During the training stage, the gating network has a tendency to reach a winner-take-all state where it always gives large weights for the same few experts. Following the previous work Shazeer et al. (2017), we add a load balancing loss function to encourage experts have equal importance. First, we count the weight of experts as loads, and calculate the square of the coefficient of variation ($\operatorname{CV}(\cdot)$) of the load values as additional loss:
$\mathcal{L}_{\text{load-balancing}}=\operatorname{CV}(\text{loads})^{2}.$ (10)
With this loss function, all experts are encouraged to have equal importance in the training stage. The total loss function is the sum of the prediction loss and load balancing loss:
$\mathcal{L}=\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{load-balancing}},$ (11)
where $\text{where }\mathcal{L}_{\text{pred}}=\operatorname{MSE}(\mathcal{Y},\hat{\mathcal{Y}}).$
3.4 Training Strategy
Although our proposed MMK is carefully designed, the performance of multiple-layer MMK is not as good as that of a single-layer MMK. From the visualization of the training loss, we infer that the main issue in training deep MMK is that the initial loss is too large and unstable. In the original KAN paper, $w_{a}$ is initialized according to the Xavier initialization and $w_{b}$ is initialized to 1 to guarantee variance-preserving Glorot and Bengio (2010); He et al. (2015); Yang and Wang (2024) with the assumption that all $x_{i}$ in $x$ are mutually independent. However, this initialization method is unstable when the number of network layers increases, because its premise, the mathematical expectation remains 0, is heavily dependent on the distribution of input data. We compare expectations of KAN and Linear in Equation 12:
$\mathbb{E}(\boldsymbol{y})^{\text{KAN}}$ $=\mathbb{E}[\phi(\boldsymbol{x})]$ (12)
$=\mathbb{E}[\boldsymbol{w}_{a}]\mathbb{E}[\mathbb{A}(\boldsymbol{x})]+\mathbb{E}[\boldsymbol{w}_{b}]\mathbb{E}[\mathbb{F}(\boldsymbol{x})],$
$=\mathbb{E}[\mathbb{F}(\boldsymbol{x})],$
$\mathbb{E}(\boldsymbol{y})^{\text{Linear}}$ $=\mathbb{E}[\boldsymbol{w}\boldsymbol{x}]=\mathbb{E}[\boldsymbol{w}]\mathbb{E}[\boldsymbol{x}]=0.$
The Linear layer keeps $\mathbb{E}=0$ by initializing $\boldsymbol{w}=0$, while KAN keeps $\mathbb{E}=0$ by supposing $\mathbb{E}[\mathbb{F}(\boldsymbol{x})]=0$. Since the real-world time series data has a biased distribution and non-negligible number of outliers, it is difficult to maintain $\mathbb{E}[\mathbb{F}(\boldsymbol{x})]=0$. Therefore, we propose a pre-sampling initialization strategy: before training starts, we forward all training samples to the model, and record the $\operatorname{Var}[\boldsymbol{x}]$ and $\operatorname{Var}[\mathbb{F}(\boldsymbol{x})]$ to initialize the $\boldsymbol{w}_{b}$ with $\mathcal{N}(0,\frac{\operatorname{Var}[\mathbb{F}(\boldsymbol{x})]}{\operatorname{Var}[\boldsymbol{x}]})$.
Dataset Variates Timesteps Granularity
ETTh1/h2 7 17,420 1 hour
ETTm1/m2 7 69,680 15 min
ECL 321 26,304 1 hour
Traffic 862 17,544 1 hour
Weather 21 52,696 10 min
Table 1: The statistics of the used datasets.
MMK iTransformer PatchTST FEDformer TimesNet SCINet TiDE DLinear
Models MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.374 0.397 0.386 0.405 0.414 0.419 0.376 0.419 0.384 0.402 0.654 0.599 0.479 0.464 0.386 0.400
192 0.419 0.429 0.441 0.436 0.460 0.445 0.420 0.448 0.436 0.429 0.719 0.631 0.525 0.492 0.437 0.432
336 0.461 0.450 0.487 0.458 0.501 0.466 0.459 0.465 0.491 0.469 0.778 0.659 0.565 0.515 0.481 0.459
ETTh1 720 0.474 0.467 0.503 0.491 0.500 0.488 0.506 0.507 0.521 0.500 0.836 0.699 0.594 0.558 0.519 0.516
96 0.301 0.353 0.297 0.349 0.302 0.348 0.358 0.397 0.340 0.374 0.707 0.621 0.400 0.440 0.333 0.387
192 0.379 0.405 0.380 0.400 0.388 0.400 0.429 0.439 0.402 0.414 0.860 0.689 0.528 0.509 0.477 0.476
336 0.432 0.446 0.428 0.432 0.426 0.433 0.496 0.487 0.452 0.452 1.000 0.744 0.643 0.571 0.594 0.541
ETTh2 720 0.446 0.463 0.427 0.445 0.431 0.446 0.463 0.474 0.462 0.468 1.249 0.838 0.874 0.679 0.831 0.657
96 0.320 0.358 0.334 0.368 0.329 0.367 0.379 0.419 0.338 0.375 0.418 0.438 0.364 0.387 0.345 0.372
192 0.364 0.383 0.377 0.391 0.367 0.385 0.426 0.441 0.374 0.387 0.439 0.450 0.398 0.404 0.380 0.389
336 0.395 0.405 0.426 0.420 0.399 0.410 0.426 0.441 0.410 0.411 0.490 0.485 0.428 0.425 0.413 0.413
ETTm1 720 0.457 0.440 0.491 0.459 0.454 0.439 0.543 0.490 0.478 0.450 0.595 0.550 0.487 0.461 0.474 0.453
96 0.176 0.261 0.180 0.264 0.175 0.259 0.203 0.287 0.187 0.267 0.286 0.377 0.207 0.305 0.193 0.292
192 0.240 0.302 0.250 0.309 0.241 0.302 0.269 0.328 0.249 0.309 0.399 0.445 0.290 0.364 0.284 0.362
336 0.299 0.342 0.311 0.348 0.305 0.343 0.325 0.366 0.321 0.351 0.637 0.591 0.377 0.422 0.369 0.427
ETTm2 720 0.397 0.401 0.412 0.407 0.402 0.400 0.421 0.415 0.408 0.403 0.960 0.735 0.558 0.524 0.554 0.522
96 0.166 0.256 0.148 0.240 0.181 0.270 0.193 0.308 0.168 0.272 0.247 0.345 0.237 0.329 0.197 0.282
192 0.187 0.274 0.162 0.253 0.188 0.274 0.201 0.315 0.184 0.289 0.257 0.355 0.236 0.330 0.196 0.285
336 0.204 0.290 0.178 0.269 0.204 0.293 0.214 0.329 0.198 0.300 0.269 0.369 0.249 0.344 0.209 0.301
ECL 720 0.247 0.323 0.225 0.317 0.246 0.324 0.246 0.355 0.220 0.320 0.299 0.390 0.284 0.373 0.245 0.333
96 0.511 0.324 0.395 0.268 0.462 0.295 0.587 0.366 0.593 0.321 0.788 0.499 0.805 0.493 0.650 0.396
192 0.529 0.330 0.417 0.276 0.466 0.296 0.604 0.373 0.617 0.336 0.789 0.505 0.756 0.474 0.598 0.370
336 0.545 0.334 0.433 0.283 0.482 0.304 0.621 0.383 0.629 0.336 0.797 0.508 0.762 0.477 0.605 0.373
Traffic 720 0.580 0.351 0.467 0.302 0.514 0.322 0.626 0.382 0.640 0.350 0.841 0.523 0.719 0.449 0.645 0.394
96 0.164 0.210 0.174 0.214 0.177 0.218 0.217 0.296 0.172 0.220 0.221 0.306 0.202 0.261 0.196 0.255
192 0.210 0.251 0.221 0.254 0.225 0.259 0.276 0.336 0.219 0.261 0.261 0.340 0.242 0.298 0.237 0.296
336 0.265 0.290 0.278 0.296 0.278 0.297 0.339 0.380 0.280 0.306 0.309 0.378 0.287 0.335 0.283 0.335
Weather 720 0.343 0.342 0.358 0.349 0.354 0.348 0.403 0.428 0.365 0.359 0.377 0.427 0.351 0.386 0.345 0.381
Rank 1 27 21 8 1 2 0 0 0
Table 2: Multivariate time series forecasting results. Forecasting horizons $P$ is 96, 192, 336, 720, and input length $T$ is set as 96. A lower value indicates better performance. The best results are marked in red.
4 Experiments
In this section, we experimentally validate the effectiveness of KAN on seven public benchmarks. Specifically, we conduct extensive experiments, including performance comparison, running speed comparison, integrating the KANs into other advanced models, and the interpretability of MMK.
with MoE w/o MoE
MoK MoL WavKAN TaylorKAN Linear
Dataset P MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.326 0.360 0.335 0.367 0.341 0.370 0.340 0.368 0.355 0.376
192 0.367 0.382 0.374 0.385 0.374 0.384 0.380 0.386 0.391 0.392
336 0.400 0.404 0.407 0.406 0.406 0.405 0.412 0.406 0.424 0.415
720 0.462 0.439 0.470 0.440 0.465 0.438 0.474 0.439 0.487 0.450
ETTm1 avg 0.389 0.396 0.397 0.400 0.397 0.399 0.402 0.400 0.414 0.407
96 0.382 0.396 0.382 0.392 0.396 0.402 0.388 0.398 0.386 0.395
192 0.430 0.426 0.436 0.423 0.439 0.430 0.438 0.424 0.437 0.424
336 0.468 0.443 0.487 0.449 0.479 0.449 0.477 0.442 0.479 0.446
720 0.450 0.458 0.486 0.472 0.464 0.461 0.477 0.461 0.481 0.470
ETTh1 avg 0.433 0.431 0.448 0.434 0.444 0.435 0.445 0.431 0.446 0.434
96 0.175 0.225 0.174 0.223 0.180 0.228 0.182 0.232 0.192 0.232
192 0.224 0.267 0.226 0.267 0.227 0.268 0.235 0.274 0.240 0.271
336 0.284 0.308 0.286 0.308 0.284 0.308 0.291 0.312 0.292 0.307
720 0.366 0.358 0.367 0.358 0.368 0.360 0.370 0.360 0.364 0.353
Weather avg 0.262 0.290 0.263 0.289 0.264 0.291 0.269 0.294 0.272 0.291
Table 3: The performance of KAN-based models vs. Linear-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
4.1 Experimental Settings
Datasets
We conduct extensive experiments on seven widely-used datasets, including ETT(h1, h2, m1, m2) Zhou et al. (2021), ECL, Traffic and Weather Lai et al. (2018), whose statistical information is shown in Table 1. We follow the same data processing operations used in TimesNet Wu et al. (2023), where the training, validation, and testing sets are divided according to chronological order.
Settings
We implement our method and baselines in a unify code library with PyTorch  [^1] https://github.com/pytorch/pytorch and Pytorch-lightning  [^2] https://github.com/Lightning-AI/pytorch-lightning , and we conduct all experiments on a GPU server with NVIDIA A100 80GB GPUs. In order to prevent the ‘Drop Last’ Qiu et al. (2024) in the test stage where the samples of the last incomplete batch are discarded, we set the test batch size to 1. We use the Mean Squared Error (MSE) and the Mean Absolute Error (MAE) as metrics. We use KAN, TaylorKAN, JacobiKAN and WavKAN to construct the MoK layer. The code in the supplementary files provides more details.
Baselines
We select seven well-known baselines, including (1) Transformer-based methods: iTransformer Liu et al. (2024a), PatchTST Nie et al. (2023) and FEDformer Zhou et al. (2022); (2) CNN-based methods: TimesNet Wu et al. (2023) and SCINet Liu et al. (2022a); (3) Linear-based methods: TiDE Das et al. and DLinear Zeng et al. (2023).
4.2 Can KANs Get the SOTA Performance?
In this section, we compare our proposed MMK with 7 popular Transformer-based, CNN-based and Linear-based baselines on seven custom datasets to comprehensively evaluate the performance of KAN. The experimental results are shown in Table 2, where the best results of all models are marked in red and the second best results are marked with underlines. The results of baselines come from previous paper Liu et al. (2024a), and the results of MMK are average of four times experiments with fixed seed [0, 1, 2, 3]. Surprisingly, MMK achieves the best performance in most cases. And it significantly outperforms on the Weather dataset, reducing the MSE by 5.74% on the 96-step forecasting compared to the second-best method, iTransformer. We infer that this improvement comes from the fact that KAN’s function-fitting ability is more suitable for processing natural science data. The performance on traffic data is poor because our method does not design a module to exploit the correlation between variables. Subsequent research can consider designing a KAN-based variable relationship mining module to better improve model performance. In general, a properly designed KAN-based model can achieve the state-of-the-art performance in most cases while maintaining the advantages of interpretability.
Dataset iTransformer +KAN +MoK
MSE MAE MSE MAE MSE MAE
ETTh1 96 0.390 0.404 0.397 0.409 0.393 0.409
192 0.445 0.436 0.443 0.436 0.442 0.441
336 0.489 0.457 0.491 0.461 0.476 0.458
720 0.505 0.487 0.510 0.488 0.494 0.485
avg 0.457 0.446 0.460 0.448 0.451 0.448
ETTm1 96 0.341 0.373 0.331 0.365 0.329 0.364
192 0.388 0.397 0.379 0.392 0.376 0.390
336 0.426 0.420 0.417 0.416 0.412 0.415
720 0.498 0.458 0.484 0.450 0.478 0.453
avg 0.413 0.412 0.403 0.406 0.399 0.406
Table 4: The performance of KAN within iTransformer. The best results of all methods are highlighted in red.
4.3 Does KANs Outperform Linear?
In this section, we conduct ablation experiments to compare KAN with Linear in time series forecasting tasks. For fair comprison, we only use one layer of forward network that means all models in this experiment only consist of RevIN and a single network layer.
  • MoK: It uses four KAN variants (KAN, WavKAN, TaylorKAN, and JacobiKAN) as experts for forecasting.
  • MoL: It uses four Linears as experts for forecasting.
  • WavKAN, TaylorKAN, Linear: They use the corresponding layer for forecasting.
The experimental results are reported in Table 3, where all the results are the average of 4 times experiments, and the best results are highlighted in red. We can conclude three useful empirical experiences. First, the KAN-based model outperforms the linear-based model in most cases. We speculate this phenomenon is due to KAN’s function representation idea, which is more efficient to capture the periodicity and trend in time series. Second, the mixture of expert structures is applicable to both KAN and Linear, which should be attributed to the fact that the gating network assigns variables to specific experts. Third, the performance of KAN-based models is affected by the specific function, which may be related to the intrinsic distribution of the time series. Through these two groups of experiments, we can conclude that KAN has competitive performance compared to Linear in time-series forecasting.
Models ETTh1 ETTm1
MSE MAE MSE MAE
TimeMixer 0.375 0.400 0.320 0.357
+MoK 0.375 0.398 0.314 0.353
Improve 0.00% 0.50% 1.88% 1.12%
GBT 0.398 0.418 0.319 0.370
+MoK 0.396 0.417 0.311 0.370
Improve 0.50% 0.24% 2.51% 0.00%
Pathformer 0.382 0.400 0.316 0.346
+MoK 0.379 0.389 0.315 0.343
Improve 0.79% 2.75% 0.32% 0.87%
Table 5: The performance of KAN within other model.
4.4 Can KANs be Integrated into Other Models?
In Section 4.2, we have demonstrated that our proposed MMK has excellent performance compared to models based on other network architectures. In order to further explore the upper bound of KAN’s performance in time series forecasting, we integrate KAN into some existing time series forecasting models in the form of replacing the original predictor so as to make full use of the progress of advanced time series forecasting technology. First, we select iTransformer Liu et al. (2024a) as baseline and replace the last linear predictor with a KAN or MoK layer. For fair comparison, we set the same network hyperparameters in all experiments where hidden dim is 512 and layer number is 2, and search the best learning rate from 1e-2 to 1e-5 through a grid searching strategy. As shown in Table 4, iTransformer with KAN and MoK can achieve better performance in most cases. Then, we intergate MoK with more models Wang et al. (2024); Chen et al. (2024); Shen et al. (2023) in Table 5, MoK can achieve better performance in most cases. These experimental findings demonstrate that MoK is a powerful form for integrating KAN into existing models.
4.5 How to Get Better MMK?
We evaluate the impact of different modules or training strategies on the forecasting performance. The pre-sampling weights initialization strategy is used in all experiments to maintain the stability of training. As shown in Figure 3, we report the MSE of $L$=96 and $P$=96, except for ‘+Longer Input’ where $L$=192 and $P$=96. First, we can find that MoK and RevIN are useful to enhance the forecasting ability. Then, adding model depth (‘+MoK’) only will drop the performance significance. The MoKBlock module and the warmup strategy can improve the performance of multi-layer MMK to achieve the best results. We also report the result under ‘+Longer Input’ to illustrate that MMK can get enhancement from longer input.
Figure 3: The impact of different modules or training strategies.
4.6 Are KANs Efficiency?
We report the size of model parameters, the training and inferring speed of KAN-based methods and baselines on ETTh1 dataset with input length 96 and prediction length 720. We set training batch size to 64, infer batch size to 1. As shown in Table 6, MMK has more parameters than DLinear, but significantly less than transformer-based methods like iTransformer and PatchTST. The running speed of MMK is slower than that of DLinear. We believe that as researchers continue to optimize the computational efficiency of KAN, the running speed of MMK will be improved in the future.
Model Param Training (it/s) Infer (it/s)
DLinear 139 K 102.51 360.93
iTransformer 3.6 M 51.03 188.79
PatchTST 7.6 M 43.92 193.58
MMK 972 K 34.86 190.57
Table 6: Comparison of model parameters and running speed of models. All results are tested with $L$=96 and $P$=720 on ETTh1.
4.7 Case Study
Figure 4: Heatmap of the expert loads in the MoK on the Weather dataset. The value at location xy equal to 1 means that the gating network assigns the y-th variable to the x-th expert in all samples.
Figure 5: The weight of each input feature.
In this section, we aim to analyze the interpretability of MMK for the time series forecasting task. First, we generate a heatmap to visualize the outputs of the gating network, which decide variable-expert assignment. The gating network outputs the matching score between variables and experts based on the input feature, and the top-$k$ matching experts are selected for each variable to predict the future state. We count the top-1 scores of all samples in the weather dataset and generate a heatmap in Figure 4, where (x, y) = 1.0 means the y-th variable matches the x-th expert in all samples. We can see that the variable is closely related to the specific expert, which indicates that MoK can effectively assign KANs to variables.
Then, we use the temperature series in the weather dataset, which collects every 10 minutes to train the KAN model. We input one daily period (the past 144 time steps) of data to predict the state of the next time step. We visualize the weight of each feature dimension of the model. As shown in Figure 5, there are three peaks in the feature weight, which are near 0, 72, and 144 time steps. 0 represents the temperature at the same time step of the previous day’s period, 144 represents the temperature at the adjacent moment, and 70 represents half a period. We conclude that KAN can learn the periodicity of time series, which can preliminary explain its effectiveness in time series prediction.
5 Conclusion
The Kolmogorov-Arnold Network (KAN) seems to provide a new perspective for interpretable time series forecasting. In this paper, we evaluate the effectiveness of KANs in time-series forecasting tasks for the first time. First, we design the mixture-of-KAN layer, which uses a mixture-of-experts structure to assign variables to best-matched KAN experts. This module aggregates the advantages of various KAN’s variants. Then, we propose the multi-layer mixture-of-KAN network (MMK), which achieves excellent performance while retaining KAN’s ability to be transformed into a combination of symbolic functions. And we use some useful tricks to address the issues in the training stage. Finally, we compare MMK with many popular baselines on seven datasets. The effectiveness of KANs in multivariate time series forecasting has been proven by extensive experiments and visualizations. This work is only a preliminary attempt to introduce KAN into time series forecasting. We hope that our work can help future studies improve the performance and interpretability of KAN-based models.